### v1

@@ Abstract
In this paper, necessary and sufficient conditions for the self-orthogonality of $t$-generator quasi-cyclic (QC) codes are presented under the Euclidean, Hermitian, and symplectic inner products, respectively. Particularly, by studying the structure of the dual codes of a class of 2-generator QC codes, we derive necessary and sufficient conditions for the QC codes to be dual-containing under the above three inner products. This class of 2-generator QC codes generalizes many known codes in the literature. Based on the above conditions, we construct several quantum stabilizer codes and quantum synchronizable codes with good parameters, some of which share parameters with certain best-known codes listed in Grassl’s code table.
@@ 1 Introduction
To ensure reliable communication, introducing error-correcting codes is an effective way to safeguard data integrity against noise and interference during transmission. Cyclic codes, an important subclass of error-correcting codes, possess efficient encoding and decoding algorithms due to their well-structured algebraic properties. As a natural generalization of cyclic codes, quasi-cyclic (QC) codes not only preserve the structural advantages of cyclic codes but also provide broader applicability and increase flexibility in parameter selection. Therefore, QC codes have attracted extensive attention and in-depth study and numerous record-breaking classical codes have been derived from QC codes Kasami1974 ; Siap2000 ; Chen2018 ; Xie2024 ; Meng2025 ; Ojiro2025 .
@@
Compared to classical communication systems, quantum communication systems are more susceptible to external disturbances. Therefore, quantum error-correcting codes (QECCs) are crucial for protecting quantum information against quantum errors, including bit-flip, phase-flip, and synchronization errors. In the 1990s, Shor et al. constructed the first QECC, marking the beginning of this field Shor1995 . Subsequently, Calderbank et al. established a fundamental connection between classical codes and quantum codes by constructing binary quantum codes from self-orthogonal classical codes Calderbank1998 . This breakthrough transforms the problem of constructing quantum codes into the task of designing classical linear codes with self-orthogonal or dual-containing properties. Since then, research on QECCs has progressed rapidly, and numerous codes with good parameters have been constructed by leveraging self-orthogonal or dual-containing classical codes Ashikhmin2001 ; Ketkar2006 ; Pang2020 ; Li2025 , most of which belong to the class of quantum stabilizer codes used to correct bit-flip and phase-flip errors. In 2013, quantum synchronizable codes (QSCs) were first proposed by Fujiwara as another subclass of QECCs Fujiwara2013 ; Fujiwara2013-2 . These codes are capable of correcting not only bit-flip and phase-flip errors but also synchronization errors. Following that, a variety of quantum synchronizable codes with desirable properties have been developed Fujiwara2014 ; Luo2018 ; Li2019 ; Luo2019 ; Dinh4ps ; Liu2021 ; Sun2023 ; Liu2023 ; Liu2024QIP . It should be emphasized that nearly all of the aforementioned QECCs are derived from classical cyclic, constacyclic, Bose–Chaudhuri–Hocquenghem (BCH) codes or algebraic geometry codes.
@@
Owing to the favorable algebraic properties of QC codes, the construction of QECCs based on QC codes has recently emerged as a prominent research focus Ling2001 . In 2018, Galindo et al. studied the dual structures of a class of 2-generator QC codes under the Euclidean, symplectic, and Hermitian inner products, and established sufficient conditions for their dual-containing property Galindo2018 . Based on these codes, they constructed quantum stabilizer codes with good parameters. Since then, significant progress has been made in the study of 1-generator and 2-generator QC codes, especially regarding their self-orthogonality and dual-containing properties under symplectic and Hermitian inner products, resulting in the construction of many high-performance quantum stabilizer codes Lv2020 ; Lv2020-2 ; Lv2021CAM ; Guan2022-2 ; Guan2022-1 ; Guan2024 ; BS2024 . To study the self-orthogonality of QC codes, it is essential to conduct a thorough analysis of the structure of their dual codes. However, the study of the dual code structure of QC codes remains challenging and scarce. As far as we know, Abdukhalikov et al. investigated the dual codes of 1-generator QC codes of index 2 KA2023 . Then Benjwal and Bhaintwal extended this result to QC codes with an arbitrary index $l$, and employed them to construct quantum stabilizer codes BS2024 . Moreover, the application of QC codes to the construction of QSCs was first realized by Du et al. in 2023, introducing new insights and methods for further development in this area duchao2023 .
@@
Inspired by the above work, we investigate the self-orthogonality property of $t$-generator QC codes with respect to the Euclidean, Hermitian and symplectic inner products, and provide the corresponding necessary and sufficient conditions for these codes to be self-orthogonal. Building on these results, we further analyze the self-orthogonal and dual-containing properties of a class of 2-generator QC codes, and derive the corresponding necessary and sufficient conditions for these properties under the three aforementioned inner products. As an application of the above analysis, we finally construct several types of quantum stabilizer codes and quantum synchronizable codes.
@@ 2 Preliminaries
Cyclic codes can be represented in multiple forms, among which the polynomial representation is the most commonly used. Denote $\mathcal{R}:=\mathbb{F}_{q}[x]/(x^{m}-1)$. There exists a bijection $\sigma$ between the codewords in $\mathcal{C}$ and the polynomials in $\mathcal{R}$, establishing a one-to-one correspondence:
@@
If $\mathcal{C}\subseteq\mathcal{C}^{\perp_{E}}$ (resp. $\mathcal{C}\subseteq\mathcal{C}^{\perp_{S}}$, $\mathcal{C}\subseteq\mathcal{C}^{\perp_{H}}$), $\mathcal{C}$ is called a Euclidean (resp. symplectic, Hermitian) self-orthogonal code. If $\mathcal{C}^{\perp_{E}}\subseteq\mathcal{C}$ (resp. $\mathcal{C}^{\perp_{S}}\subseteq\mathcal{C}$, $\mathcal{C}^{\perp_{H}}\subseteq\mathcal{C}$), $\mathcal{C}$ is called a Euclidean (resp. symplectic, Hermitian) dual-containing code. It is easy to see that $\mathcal{C}$ is dual-containing if and only if its dual code is self-orthogonal under any type of inner products.
To discuss the properties of self-orthogonality or dual-containing of QC codes, we present some simple notations.
@@
$k^{[q]}(x)$ $=k_{0}^{q}+k_{1}^{q}x+\cdots+k_{m-1}^{q}x^{m-1},$
$k^{*}(x)$ $=x^{\deg(k(x))}k(x^{-1})\pmod{x^{m}-1},$
$\overline{k}(x)$ $=k_{0}+k_{m-1}x+k_{m-2}x^{2}+\cdots+k_{1}x^{m-1}=k(x^{-1})\pmod{x^{m}-1},$
$k^{\perp}(x)$ $=f(0)^{-1}{f^{*}(x)},f(x)=\frac{x^{m}-1}{\gcd(k(x),x^{m}-1)},$
$k^{\perp_{H}}(x)$ $=(k^{[q]}(x))^{\perp}.$
@@
    $\overline{\overline{k_{1}}}(x)=k_{1}(x)$, $\overline{k_{1}(x)k_{2}(x)}=\overline{k_{1}}(x)\overline{k_{2}}(x)$ and $\overline{k_{1}(x)+k_{2}(x)}=\overline{k_{1}}(x)+\overline{k_{2}}(x)$ for any $k_{1}(x),k_{2}(x)\in\mathcal{R}$;
@@
Similarly to cyclic codes, QC codes can also be represented in polynomial form. Let $\mathcal{C}$ be a QC code of length $ml$ and index $l$ over $\mathbb{F}_{q}$. For any codeword $\mathbf{c}=(c_{0,0},\ldots,c_{l-1,0},c_{0,1},\ldots,c_{l-1,1},\ldots,c_{0,m-1},\ldots,c_{l-1,m-1})\in\mathcal{C}$, we define a map as follows:
@@
Ling2001 For $\mathbf{u},\mathbf{v}\in\mathbb{F}_{q}^{ml}$, the Euclidean inner product of $\mathbf{u}$ and its all $l$ right cyclic shifts with $\mathbf{v}$ is zero if and only if $\phi(\mathbf{u})*\phi(\mathbf{v})=\sum\limits_{i=0}^{l-1}u_{i}(x)\overline{v_{i}}(x)=0$.
For the polynomial $k(x)=k_{0}+k_{1}x+k_{2}x^{2}+\dots+k_{m-1}x^{m-1}\in\mathcal{R}$, denote $[k(x)]=(k_{0},k_{1},\dots,k_{m-1})$. Then $k(x)$ uniquely determines the following circulant matrix
@@
where each $K_{ij}$ is a circulant matrix corresponding to a polynomial $k_{ij}(x)$ over $\mathcal{R}$, $1\leq i\leq t$ and $1\leq j\leq l$. A QC code whose generator matrix has the form $G$ is called a $t$-generator QC code with index $l$. And the generator set can be expressed as $\{\vec{K}_{1},\vec{K}_{2},\dots,\vec{K}_{t}\}$, where $\vec{K}_{i}=([k_{i1}(x)],[k_{i2}(x)],\dots,[k_{il}(x)])$, $1\leq i\leq t$.
Let $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ be two $t$-generator QC codes of length $ml$ and index $l$ over $\mathbb{F}_{q}$ with the generator sets $\{\vec{a}_{1},\vec{a}_{2},\dots,\vec{a}_{t}\}$ and $\{\vec{b}_{1},\vec{b}_{2},\dots,\vec{b}_{t}\}$, respectively. If $\vec{b}_{1},\vec{b}_{2},\dots,\vec{b}_{t}$ can be linearly represented by $\vec{a}_{1},\vec{a}_{2},\dots,\vec{a}_{t}$, we have $\mathcal{C}_{2}\subseteq\mathcal{C}_{1}$.
@@
Seguin2004 Let $\mathcal{C}$ be a 1-generator QC code of length $ml$ and index $l$ over $\mathbb{F}_{q}$, of the form
@@
where $\mathcal{R}=\mathbb{F}_{q}[x]/\langle x^{m}-1\rangle$. Then $g(x)=\gcd\left(a_{1}(x),a_{2}(x),\dots,a_{l}(x),x^{m}-1\right)$, $h(x)=\frac{x^{m}-1}{g(x)}$ are called the generator polynomial and parity-check polynomial of $\mathcal{C}$ respectively. Further, the dimension of $\mathcal{C}$ is given by $\dim(\mathcal{C})=m-\deg(g(x))=\deg(h(x))$.
An $[[n,k,d]]_{q}$ quantum stabilizer code is a $q$-ary quantum code that encodes $k$ logical qudits into $n$ physical qudits and can detect all errors of weight less than $d$, where $q$ is a power of a prime number. Such a code is defined as the joint eigenspace of an abelian subgroup of the error group acting on the Hilbert space $\mathbb{C}^{q^{n}}$.
An $(a_{l},a_{r})$-$[[n,k]]_{q}$ quantum synchronizable code is a $q$-ary quantum code that encodes $k$ qudits into $n$ physical qudits and is capable of correcting misalignment of up to $a_{l}$ qudits to the left and $a_{r}$ qudits to the right.
@@ 3 Necessary and Sufficient Conditions of $t$-Generator QC Codes to be Self-Orthogonal
$\mathcal{C}$ is a $t$-generator QC code of index $l$ over $\mathbb{F}_{q}$, generated by
@@
The code $\mathcal{C}$ is Euclidean self-orthogonal if and only if
$h_{r}(x)\mid\overline{g_{s}}(x)\sum_{j=1}^{l}k_{rj}(x)\overline{k_{sj}}(x),$
where $1\leq r\leq s\leq t$.
@@
$([x^{j_{s}}k_{s1}(x)g_{s}(x)],[x^{j_{s}}k_{s2}(x)g_{s}(x)],\cdots,[x^{j_{s}}k_{sl}(x)g_{s}(x)])\rangle_{E}=0,$
for all $10\leq i_{r}\leq m-\deg(g_{r}(x))-$ 1, $10\leq j_{s}\leq m-\deg(g_{s}(x))-$ 1, $1\leq r,s\leq t$. By Lemma 2, the above equation holds if and only if $x^{i_{r}+m-j_{s}}g_{r}(x)\overline{g_{s}}(x)\sum_{u=1}^{l}k_{ru}(x)\overline{k_{su}}(x)\equiv 0\mod{x^{m}-1}$ i.e. $h_{r}(x)\mid\overline{g_{s}}(x)\sum_{u=1}^{l}k_{ru}(x)\overline{k_{su}}(x)$, for any $r,s\in\{1,2,\cdots,t\}$. Due to the symmetry of the Euclidean inner product, it is sufficient to restrict our attention to the case of $1\leq r\leq s\leq t$. This completes the proof. ∎
Example 1.
Let $q=2$ and $m=21$. Choose $g(x)=x^{10}+x^{8}+x^{6}+x^{4}+x^{3}+1$, $v_{1}(x)=x^{3}+x+1$, and $v_{2}(x)=x^{3}+x^{2}$. Then we can construct a 1-generator QC code $\mathcal{C}=\langle([v_{1}(x)g(x)],\ [v_{2}(x)g(x))]\rangle$. Using Magma for computation, we find that $\mathcal{C}$ is a binary linear code with parameters $[42,11,16]_{2}$, which is a best-known linear code (BKLC); see codetables . According to Theorem 1, the code $\mathcal{C}$ is Euclidean self-orthogonal.
Remark 1.
(1) It should be noted that the above result can be viewed as a generalization of the conclusion in BS2024 . Specifically, Theorem 1 reduces to Theorem 11 in BS2024 when $t=1$.
(2) Under the Hermitian inner product, similar necessary and sufficient conditions for the self-orthogonality of QC codes can also be established. The only modification required is to replace $GG^{T}=0$ with $G^{[q]}G^{T}=0$, where $G^{[q]}$ denotes the matrix obtained by raising each entry of $G$ to the $q$-th power. In the following, we present the corresponding result without providing the proof.
Theorem 2.
Let $\mathcal{C}$ be the code given in Definition 3, except that the base field is replaced by $\mathbb{F}_{q^{2}}$. Then, $\mathcal{C}$ is Hermitian self-orthogonal if and only if
$h_{r}^{[q]}(x)\mid\overline{g_{s}}(x)\sum_{j=1}^{l}k_{rj}^{[q]}(x)\overline{k_{sj}}(x),$
where $1\leq r\leq s\leq t$.
To establish necessary and sufficient conditions for symplectic self-orthogonality of QC codes, we need the following lemma.
Lemma 3.
Galindo2018 Let $f(x)$, $g(x)$, and $k(x)$ be polynomials over $\mathcal{R}$. Then, the following property holds for the Euclidean inner product:
Theorem 3.
The code $\mathcal{C}$ is symplectic self-orthogonal if and only if
$h_{r}(x)\mid\overline{g_{s}}(x)\sum_{j=1}^{\omega}(k_{rj}(x)\overline{k_{s(\omega+j)}}(x)-k_{r(\omega+j)}(x)\overline{k_{sj}}(x)),$ (1)
where $1\leq r\leq s\leq t$, $l$ is even and $\omega=\frac{l}{2}$.
Proof.
First of all, it should be noted that $l$ must be even under the symplectic inner product. Furthermore, for the sake of simplicity, we abbreviate polynomial expressions, such as writing $g_{r}(x)$ simply as $g_{r}$, provided that no ambiguity is caused.
From the proof of Theorem 1, we have determined the structure of the generator matrix of $\mathcal{C}$. Then $\mathcal{C}$ is symplectic self-orthogonal if and only if the symplectic inner product of any two rows of the generator matrix is zero, that is
@@
for all $10\leq i_{r}\leq m-\deg(g_{r}(x))-$ 1, $10\leq j_{s}\leq m-\deg(g_{s}(x))-$ 1, and $1\leq r,s\leq t$. This is equivalent to
$,\displaystyle\langle([x^{i_{r}}k_{r1}g_{r}],$$\cdots,[x^{i_{r}}k_{r\omega}g_{r}]),([x^{j_{s}}k_{s(\omega+1)}g_{s}],\cdots,[x^{j_{s}}k_{sl}g_{s}])\rangle_{E}$
$-\langle([x^{i_{r}}k_{r(\omega+1)}g_{r}],\cdots,[x^{i_{r}}k_{rl}g_{r}]),([x^{j_{s}}k_{s1}g_{s}],\cdots,[x^{j_{s}}k_{s\omega}g_{s}])\rangle_{E}=0.$ (2)
According to Lemma 3, Eq. (2) can be transformed into

### v2

@@ Abstract
In this paper, necessary and sufficient conditions for the self-orthogonality of $t$-generator quasi-cyclic (QC) codes are presented under the Euclidean, Hermitian, and symplectic inner products, respectively. Particularly, by studying the structure of the dual codes of a class of 2-generator QC codes, we derive necessary and sufficient conditions for the QC codes to be dual-containing under the above three inner products. This class of 2-generator QC codes generalizes many known codes in the literature. Based on the above conditions, we construct several quantum stabilizer codes and quantum synchronizable codes with good parameters, some of which share parameters with certain best-known codes listed in the codetable of Grassl.
@@ 1 Introduction
To ensure the reliability of classical communication, introducing error-correcting codes is an effective way to safeguard data integrity against noise and interference during transmission. Cyclic codes, an important subclass of error-correcting codes, possess efficient encoding and decoding algorithms due to their well-structured algebraic properties. As a natural generalization of cyclic codes, quasi-cyclic (QC) codes not only preserve the structural advantages of cyclic codes but also provide broader applicability and increase flexibility in parameter selection. Therefore, QC codes have attracted extensive attention and in-depth study and numerous record-breaking classical codes have been derived from QC codes [1, 2, 3, 4, 5].
@@
Compared to classical communication systems, quantum communication systems are more susceptible to external disturbances. Therefore, quantum error-correcting codes (QECCs) are crucial for protecting quantum information against quantum errors, including bit-flip, phase-flip, and synchronization errors. In the 1990s, Shor et al. constructed the first QECC, marking the beginning of this field [6]. Subsequently, Calderbank et al. showed that binary quantum codes can be systematically derived from self‑orthogonal classical codes, thereby recasting the construction of quantum codes as the design of classical linear codes with self‑orthogonal or dual‑containing properties [7].
Since then, research on QECCs has progressed rapidly, and numerous codes with good parameters have been constructed by leveraging self-orthogonal or dual-containing classical codes [8, 9, 10, 11], most of which belong to the class of quantum stabilizer codes used to correct bit-flip and phase-flip errors. In 2013, quantum synchronizable codes (QSCs) were first proposed by Fujiwara as another subclass of QECCs [12, 13]. These codes are capable of correcting not only bit-flip and phase-flip errors but also synchronization errors. Following that, a variety of quantum synchronizable codes with desirable properties have been developed [14, 15, 16, 17, 18, 19, 20, 21, 22]. It should be emphasized that nearly all of the aforementioned QECCs are derived from classical cyclic, constacyclic, Bose–Chaudhuri–Hocquenghem (BCH) codes or algebraic geometry codes.
@@
Owing to the favorable algebraic properties of QC codes, the construction of QECCs based on QC codes has recently emerged as a prominent research focus [23]. In 2018, Galindo et al. studied the dual structure of a class of 2-generator QC codes under the Euclidean, symplectic, and Hermitian inner products, and established sufficient conditions for their dual-containing property. Based on these codes, they constructed quantum stabilizer codes with good parameters [24]. Since then, significant progress has been made in the study of 1-generator and 2-generator QC codes, especially regarding their self-orthogonality and dual-containing properties under symplectic and Hermitian inner products, resulting in the construction of many high-performance quantum stabilizer codes [25, 26, 27, 28, 29, 30, 31]. In addition, in order to obtain quantum stabilizer codes with better performance, Ezerman et al. also established the necessary and sufficient conditions for the self-orthogonality of certain quasi-twisted codes from the perspective of direct sum decomposition [32].
To investigate the dual-containing properties of QC codes, it is essential to conduct a thorough analysis of the structure of their dual codes. However, the study of the dual code structure of QC codes remains challenging and scarce. As far as we know, Abdukhalikov et al. investigated the dual codes of 1-generator QC codes of index 2 [33]. Then Benjwal and Bhaintwal extended this result to QC codes with an arbitrary index $l$, and employed them to construct quantum stabilizer codes [31]. Moreover, the application of QC codes to the construction of QSCs was first realized by Du et al. in 2023, introducing new insights and methods for further development in this area [34].
@@
Inspired by the above work, we investigate the self-orthogonality of $t$-generator QC codes with respect to the Euclidean, Hermitian and symplectic inner products, and provide the corresponding necessary and sufficient conditions for these codes to be self-orthogonal. Building upon the above results, we further investigate the algebraic structure of the duals of a class of 2-generator QC codes and derive the necessary and sufficient conditions for them to be self-orthogonal and dual-containing under the three aforementioned inner products. As an application of the above analysis, we finally construct several types of quantum stabilizer codes and quantum synchronizable codes.
@@ 2 Preliminaries
Cyclic codes can be represented in various forms. The polynomial representation is the most commonly used. Denote $\mathcal{R}:=\mathbb{F}_{q}[x]/(x^{m}-1)$. There exists a bijection $\sigma$ between the codewords in $\mathcal{C}$ and the polynomials in $\mathcal{R}$, establishing a one-to-one correspondence:
@@
If $\mathcal{C}\subseteq\mathcal{C}^{\perp_{E}}$ (resp. $\mathcal{C}\subseteq\mathcal{C}^{\perp_{S}}$, $\mathcal{C}\subseteq\mathcal{C}^{\perp_{H}}$), $\mathcal{C}$ is called a Euclidean (resp. symplectic, Hermitian) self-orthogonal code. If $\mathcal{C}^{\perp_{E}}\subseteq\mathcal{C}$ (resp. $\mathcal{C}^{\perp_{S}}\subseteq\mathcal{C}$, $\mathcal{C}^{\perp_{H}}\subseteq\mathcal{C}$), $\mathcal{C}$ is called a Euclidean (resp. symplectic, Hermitian) dual-containing code. It is easy to see that $\mathcal{C}$ is dual-containing if and only if its dual code is self-orthogonal under the corresponding type of inner products.
To facilitate the discussion of self-orthogonality and dual-containing properties of QC codes, we begin by introducing some notations.
@@
$k^{[q]}(x)$ $=k_{0}^{q}+k_{1}^{q}x+\cdots+k_{m-1}^{q}x^{m-1};\quad\quad k^{*}(x)=x^{\deg(k(x))}k(x^{-1})\pmod{x^{m}-1}$;
$\overline{k}(x)$ $=k_{0}+k_{m-1}x+k_{m-2}x^{2}+\cdots+k_{1}x^{m-1}=k(x^{-1})\pmod{x^{m}-1}$;
$k^{\perp}(x)$ $=f(0)^{-1}{f^{*}(x)},f(x)=\frac{x^{m}-1}{\gcd(k(x),x^{m}-1)};\quad\quad k^{\perp_{H}}(x)=(k^{[q]}(x))^{\perp}.$
@@
    $\overline{\overline{k_{1}}}(x)=k_{1}(x)$, $\overline{k_{1}(x)k_{2}(x)}=\overline{k_{1}}(x)\overline{k_{2}}(x)$, and $\overline{k_{1}(x)+k_{2}(x)}=\overline{k_{1}}(x)+\overline{k_{2}}(x)$ for any $k_{1}(x),k_{2}(x)\in\mathcal{R}$;
@@
Proof.
The verification of the first five statements is straightforward and thus omitted. We provide a detailed proof only for the final one. Our primary objective is to prove the following equation
$(k^{[q]}(x))^{\perp}=\big{(}k^{\perp}(x)\big{)}^{[q]}.$
Recall that
$k^{\perp}(x)=f(0)^{-1}f^{*}(x),\quad\text{where}\quad f(x)=\frac{x^{m}-1}{\gcd(k(x),x^{m}-1)}.$
Then,
$\big{(}k^{\perp}(x)\big{)}^{[q]}=f(0)^{-q}(f^{*}(x))^{[q]}.$
Note that
$\frac{x^{m}-1}{\gcd(k^{[q]}(x),x^{m}-1)}=\frac{x^{m}-1}{\gcd(k(x),x^{m}-1)^{[q]}}=\left(\frac{x^{m}-1}{\gcd(k(x),x^{m}-1)}\right)^{[q]}=f^{[q]}(x)$
and $(f^{*}(x))^{[q]}=\big{(}f^{[q]}(x)\big{)}^{*}.$ Also, $f^{[q]}(0)=f(0)^{q}$. Thus
$f(0)^{-q}(f^{[q]}(x))^{*}=f^{[q]}(0)^{-1}(f^{[q]}(x))^{*}=(k^{[q]}(x))^{\perp}.$
Therefore,
$(k^{[q]}(x))^{\perp}=\big{(}k^{\perp}(x)\big{)}^{[q]},$
which proves the result. ∎
As with cyclic codes, QC codes can also be expressed in polynomial form. Let $\mathcal{C}$ be a QC code of length $ml$ and index $l$ over $\mathbb{F}_{q}$. For any codeword $\mathbf{c}=(c_{0,0},\ldots,c_{l-1,0},c_{0,1},\ldots,c_{l-1,1},\ldots,c_{0,m-1},\ldots,c_{l-1,m-1})\in\mathcal{C}$, a map is defined as follows:
@@
[23] For any $\mathbf{u},\mathbf{v}\in\mathbb{F}_{q}^{ml}$, the Euclidean inner product of $\mathbf{v}$ with $\mathbf{u}$ and its all $l$ right cyclic shifts is zero if and only if $\phi(\mathbf{u})*\phi(\mathbf{v})=\sum\limits_{i=0}^{l-1}u_{i}(x)\overline{v_{i}}(x)=0$.
For the polynomial $k(x)=k_{0}+k_{1}x+k_{2}x^{2}+\dots+k_{m-1}x^{m-1}\in\mathcal{R}$, denote $k(x)]=(k_{0},k_{1},\dots,k_{m-1})$. Then $k(x)$ uniquely determines the following circulant matrix
@@
where each $K_{ij}$ is a circulant matrix corresponding to a polynomial $k_{ij}(x)$ in $\mathcal{R}$, $1\leq i\leq t$ and $1\leq j\leq l$. A QC code whose generator matrix has the form $G$ is called a $t$-generator QC code with index $l$ and its generating set can be expressed as $\{\vec{K}_{1},\vec{K}_{2},\dots,\vec{K}_{t}\}$, where $\vec{K}_{i}=([k_{i1}(x)],[k_{i2}(x)],\dots,[k_{il}(x)])$, $1\leq i\leq t$.
Let $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ be two $t$-generator QC codes of length $ml$ and index $l$ over $\mathbb{F}_{q}$ with the generating sets $\{\vec{a}_{1},\vec{a}_{2},\dots,\vec{a}_{t}\}$ and $\{\vec{b}_{1},\vec{b}_{2},\dots,\vec{b}_{t}\}$, respectively. If $\vec{b}_{1},\vec{b}_{2},\dots,\vec{b}_{t}$ can be linearly represented by $\vec{a}_{1},\vec{a}_{2},\dots,\vec{a}_{t}$, we have $\mathcal{C}_{2}\subseteq\mathcal{C}_{1}$.
@@
[35] Let $\mathcal{C}$ be a 1-generator QC code of length $ml$ and index $l$ over $\mathbb{F}_{q}$, of the form
@@
where $\mathcal{R}=\mathbb{F}_{q}[x]/(x^{m}-1)$. Then $g(x)=\gcd\left(a_{1}(x),a_{2}(x),\dots,a_{l}(x),x^{m}-1\right)$, $h(x)=\frac{x^{m}-1}{g(x)}$ are called the generator polynomial and parity-check polynomial of $\mathcal{C}$ respectively. Further, the dimension of $\mathcal{C}$ is given by $\dim(\mathcal{C})=m-\deg(g(x))=\deg(h(x))$.
An $n,k,d]]_{q}$ quantum stabilizer code is a $q$-ary quantum code that encodes $k$ logical qudits into $n$ physical qudits and can detect all errors of weight less than $d$, where $q$ is a power of a prime number. Such a code is defined as the joint eigenspace of an abelian subgroup of the error group acting on the Hilbert space $\mathbb{C}^{q^{n}}$.
An $(a_{l},a_{r})$-$n,k]]_{q}$ quantum synchronizable code is a $q$-ary quantum code that encodes $k$ qudits into $n$ physical qudits and is capable of correcting misalignment of up to $a_{l}$ qudits to the left and $a_{r}$ qudits to the right.
@@ 3 Necessary and sufficient conditions of $t$-generator QC codes to be self-orthogonal
$\mathcal{C}$ is a $t$-generator QC code of index $l$ over $\mathbb{F}_{q}$ generated by
@@
Lemma 3.
[24] Let $f(x)$, $g(x)$, and $k(x)$ be polynomials over $\mathcal{R}$. Then, the following property holds for the Euclidean inner product:
@@
Let $\mathcal{C}$ be the code in Definition 3. Then the necessary and sufficient conditions for $\mathcal{C}$ to be self-orthogonal with respect to the following inner products are given below:
  1. Case A.
    $\mathcal{C}$ is Euclidean self-orthogonal if and only if
    $h_{r}(x)\mid\overline{g_{s}}(x)\sum_{j=1}^{l}k_{rj}(x)\overline{k_{sj}}(x),\quad\text{where}\quad 1\leq r\leq s\leq t;$
  2. Case B.
    $\mathcal{C}$ is Hermitian self-orthogonal if and only if
    $h_{r}^{[q]}(x)\mid\overline{g_{s}}(x)\sum_{j=1}^{l}k_{rj}^{[q]}(x)\overline{k_{sj}}(x),\quad\text{where}\quad 1\leq r\leq s\leq t;$
  3. Case C.
    $\mathcal{C}$ is Symplectic self-orthogonal if and only if
    $h_{r}(x)\mid\overline{g_{s}}(x)\sum_{j=1}^{\omega}(k_{rj}(x)\overline{k_{s(\omega+j)}}(x)-k_{r(\omega+j)}(x)\overline{k_{sj}}(x)),$
    where $1\leq r\leq s\leq t$, $l$ is even and $\omega=\frac{l}{2}$.
@@
$([x^{j_{s}}k_{s1}(x)g_{s}(x)],[x^{j_{s}}k_{s2}(x)g_{s}(x)],\cdots,[x^{j_{s}}k_{sl}(x)g_{s}(x)])\rangle_{E}=0,$ (3.1)
for all $0\leq i_{r}\leq m-\deg(g_{r}(x))-1$, $0\leq j_{s}\leq m-\deg(g_{s}(x))-1$, $1\leq r,s\leq t$. By Lemma 2, Eq. (3.1) holds if and only if
$x^{i_{r}+m-j_{s}}g_{r}(x)\overline{g_{s}}(x)\sum_{u=1}^{l}k_{ru}(x)\overline{k_{su}}(x)\equiv 0\mod{x^{m}-1},$
i.e., $h_{r}(x)\mid\overline{g_{s}}(x)\sum_{u=1}^{l}k_{ru}(x)\overline{k_{su}}(x)$, for any $r,s\in\{1,2,\cdots,t\}$. Due to the symmetry of the Euclidean inner product, it is sufficient to restrict our attention to the case of $1\leq r\leq s\leq t$.
Under the Hermitian inner product, similar necessary and sufficient conditions for the self-orthogonality of QC codes can also be established. The only modification required is to replace $GG^{T}=0$ with $G^{[q]}G^{T}=0$, where $G^{[q]}$ denotes the matrix obtained by raising each entry of $G$ to the $q$-th power. We omit the proof of this part here.
For notational convenience, polynomial expressions like $g_{r}(x)$ are abbreviated as $g_{r}$, as long as the meaning remains unambiguous. Then $\mathcal{C}$ is symplectic self-orthogonal if and only if the symplectic inner product of any two rows of the generator matrix is zero, that is,
@@
for all $0\leq i_{r}\leq m-\deg(g_{r}(x))-1$, $0\leq j_{s}\leq m-\deg(g_{s}(x))-1$, and $1\leq r,s\leq t$. This is equivalent to
$\langle([x^{i_{r}}$ $k_{r1}g_{r}],\cdots,[x^{i_{r}}k_{r\omega}g_{r}]),([x^{j_{s}}k_{s(\omega+1)}g_{s}],\cdots,[x^{j_{s}}k_{sl}g_{s}])\rangle_{E}$
$-\langle([x^{i_{r}}k_{r(\omega+1)}g_{r}],\cdots,[x^{i_{r}}k_{rl}g_{r}]),([x^{j_{s}}k_{s1}g_{s}],\cdots,[x^{j_{s}}k_{s\omega}g_{s}])\rangle_{E}=0.$ (3.2)
According to Lemma 3, Eq. (3.2) can be transformed into